### v1

# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation

Abstract
Phishing email is a serious cyber threat that tries to deceive users by sending false emails with the intention of stealing confidential information or causing financial harm. Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make detection and prevention of phishing more challenging. Despite extensive academic research, phishing detection remains an ongoing and formidable challenge in the cybersecurity landscape. Large Language Models (LLMs) and Masked Language Models (MLMs) possess immense potential to offer innovative solutions to address long-standing challenges. In this research paper, we present an optimized, fine-tuned transformer-based DistilBERT model designed for the detection of phishing emails. In the detection process, we work with a phishing email dataset and utilize the preprocessing techniques to clean and solve the imbalance class issues. Through our experiments, we found that our model effectively achieves high accuracy, demonstrating its capability to perform well. Finally, we demonstrate our fine-tuned model using Explainable-AI (XAI) techniques such as Local Interpretable Model-Agnostic Explanations (LIME) and Transformer Interpret to explain how our model makes predictions in the context of text classification for phishing emails.
keywords:
[…]
1 Introduction
The rapid advancement of internet technology has significantly changed the way people communicate online while also raising more security issues (Salloum et al., 2021). One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation. In this current decade, email phishing has changed throughout time to become a more effective method for carrying out fraudulent actions. As phishers continuously innovate new techniques, detecting phishing has become increasingly challenging, posing a formidable task for law enforcement agencies, governments, and organizations combating this cyber threat. To evade falling victim to this threat, organizations need to adeptly detect and prevent these attacks at scale by adopting effective phishing detection techniques. Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023). Moreover, because of the rapid advancements in technology, new methods based on AI are being developed daily in the cybersecurity field.
Transformer-based models become more popular in the context of text classification. The transformer is an innovative neural design that uses the attention mechanism to encode incoming data into powerful characteristics (Han et al., 2021). Transformer-based models have greatly improved the processing, comprehension, and interpretation of text data inputs, which has had a revolutionary effect on the development of phishing classification techniques (Jamal et al., 2023). Large Language Models like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023). The increasing complexity of LLMs is radically changing artificial intelligence and has the potential to transform algorithm development and application by understanding human language through extensive text data patterns and correlations.
Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as Large Language Models because of their ability to comprehend and produce text that is similar to that of a human. BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to grasp contextual word embeddings (Koroteev, 2021), become increasingly efficient and valuable in text classification and generation, including tasks like spam and phishing detection, sentiment analysis (Singh et al., 2021), and more other fields. BERT is designed to take into account both left and right contexts at every level to pre-train complex bidirectional representations from unlabeled text (Devlin et al., 2018). With simply one extra output layer, this method allows the pre-trained BERT model to be refined, leading to the development of state-of-the-art models for a variety of applications. In order to model language, BERT uses a Masked Language Model, in which certain words in a sentence are randomly replaced with a unique token (usually [MASK]). Next, by examining the contextual clues offered by nearby words, the model is trained to infer the original words. BERT makes use of a Transformer architecture technique called “self-attention”, which allows it to recognize inter-word relationships in text sequences and assess the relative importance of individual words within a sentence in order to understand the context. Based on the transformer layers number and parameters BERT model has different variants such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
Despite BERT’s versatility across NLP tasks, comprehending its internal mechanisms often poses challenges, resembling a “black box” due to its complexity. Unraveling how the model operates and forms predictions demands considerable effort (Xu et al., 2019). XAI techniques emerge as a vital solution, offering significant perceptions of the model’s reasoning and illuminating the decision-making process (Hoffman et al., 2018; Anan et al., 2023). Recently, Sarker (Sarker, 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested to explore more with experiments. XAI techniques can illuminate the words or features deemed most influential by the BERT model in specific text classification tasks of any nature. In this research, our main goal is to show transparency in phishing email detection with explainable AI. The major contribution of the work is as follows:
  • We present a transformer-based Phishing email detection model, a fine-tuned version of DistilBERT.
  • Two different XAI techniques LIME and Transformer Interpret are utilized to explore the interpretability of the proposed fine-tuned model and visualize the scoring percentage of those techniques on prediction.
  • We demonstrate our model performance on both imbalanced and balanced datasets.
The remaining sections of the paper are structured as follows: Section 2 delves into related research works. Section 3 offers a comprehensive explanation of the proposed model’s framework and methodology. In Section 4, the experimental outcomes and results are elaborated upon. Section 5 examines the transparency of the model using XAI. Moreover, within Section 6, there is an extensive discussion of the results. Lastly, Section 7 presents the concluding remarks and outlines the future prospects of this work.
2 Literature Review
The field of cybersecurity has grown increasingly important and complex in the modern world. There are several subfields within cybersecurity, including data security, application security, network security, etc. Cyber-attacks of all kinds are becoming more and more frequent, and they can destroy or steal imported data. In addition, in this cybersecurity field, there are an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks. For the improvement of phishing email detection and prevention systems, many researchers have developed a wide range of novel strategies to counteract phishing emails, utilizing state-of-the-art methods from deep learning, machine learning, and hybrid models. Phishing is still a serious cybersecurity concern and one of the main ways hackers breach networks.
Detecting phishing attempts early on is critical for businesses nowadays. The security of an organization can be significantly increased by machine learning, which can identify potential assaults early on (Apruzzese et al., 2023). In ref. (Yasin and Abuhasan, 2016), a text processing, data mining, and knowledge discovery approach was combined to create an intelligent phishing email detection model. They utilized text stemming and WordNet ontology in pre-processing to enrich the model with synonyms and employed five machine learning classification algorithms (J48, Naïve Bayes, Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Random Forest (RF)). They have notably achieved 99.1% accuracy with RF and 98.4% using J48 for an accredited dataset. The study (Harikrishnan et al., 2018) employed techniques such as SVD and NMF for feature extraction and dimensionality reduction from legitimate and phishing email text data. For the numeric representation of phishing emails, the author integrated TFIDF with the SVD and NMF, and Decision Trees, Logistic Regression, Random Forest, Naive Bayes, K-Nearest Neighbors (KNN), AdaBoost, and SVM have been utilized as classifiers. Hamid et al. (2013) presented a hybrid feature selection technique combining content-based and behaviour-based approaches to detect phishing emails. Identifying the behaviour-based features in phishing emails was the main objective of this research because an attacker cannot disguise these types of features. They examined the overall process with a phishing dataset that consists of 6923 emails where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset and achieved 94% accuracy with this hybrid feature selection technique. With an emphasis on the importance of top-ranking features, the study (Zamir et al., 2020) examines feature selection methods (GR, IG, Relief-F, RFE) on a dataset of 11,055 phishing sites with 32 attributes. To improve classifier accuracy, it suggests new attributes that are generated from the strongest and weakest features. A variety of supervised learning methods with differing feature sets are compared, such as RF, SVM, bagging, kNN, Neural Network (NN), and j48. The proposed method using RF showed 97.3% and using stacking achieved 97.4% accuracy.
In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts.Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP). A publicly available balanced phishing dataset has been utilized where the authors used the email body’s text to extract features that improved the proposed model’s accuracy. The model showed 98.2% accuracy with a low false-positive rate of 0.015%. For phishing detection, a new Intelligent Cuckoo Search (CS) Optimization Algorithm with a Deep Learning-based Phishing Email Detection and Classification (ICSOA-DLPEC) approach has been proposed in (Brindha et al., 2023). Data preprocessing (email cleaning, tokenization, and stop-word elimination) techniques were employed to derive the important features from the Enron email dataset. The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameter of the proposed model and achieved a high accuracy of 99.72%.Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam email detection. They developed a phish responder which is a Python-based command line solution and used deep learning and NLP for this classification task. The proposed LSTM model achieved 99% accuracy for the text-based dataset. Fang et al. (2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural networks (RCNN) model which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM). 99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced where the positive rate (FPR) is 0.043%. Other deep learning algorithms, such the reinforcement learning, Recurrent Neural Network (RNN), and Deep Belief Network (DBN), are also in use (Zhang and Li, 2017; Bahnsen et al., 2017; Smadi et al., 2018).
Moreover, transformer-based models are also making a good impact in the cybersecurity field.Thapa et al. (2023) concentrated on two specific models: RNN and BERT to test how well these models could identify phishing emails. They looked at how well these models performed when the email data was distributed unevenly across several organizations and when added more data. In this experiment, the BERT model’s test accuracy was 96.1%. In addition, another study (Atawneh and Aljehani, 2023) focused on different deep learning architectures (CNNs, LSTMs, RNNs, BERT) to identify the emails as legitimate or phishing by analyzing different features of emails. The proposed model using BERT and LSTM, reaching an impressive accuracy of 99.61% on publicly available 4 different datasets, outperforming other methods. Nowadays, other variants of the BERT model such as DistilBERT, RoBERTa, and TinyBER (Jiao et al., 2019) are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023). The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants such as the fine-tuned DistilBERT and RoBERTA model and showed good accuracy in this research scope. Lee et al. Lee et al. (2020) proposed Context-Aware Tiny Bert (CatBERT) which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos. The detection rate of this proposed model was 87% which was far better than other compared models. Furthermore, researchers also used fine-tuned BERT-based models for phishing URL detection (Wang et al., 2023; Maneriker et al., 2021).
Overall, this literature survey shows the existing works although showing good performance by using different datasets and models but most of the research didn’t work model’s interpretability or explainability. “How does a model work?” or “Which things or materials contribute to detecting phishing?” those types of questions can arise in their working procedure. In this work, to answer those types of questions and solve the existing drawbacks in the literature, we have worked with explainable AI using not only one technique but also used two different techniques with different methodologies. In the literature, although some researchers did full experiments for phishing email detection but they utilized spam datasets as phishing datasets. In cybersecurity, we know that two different types of emails exist: spam and phishing, and not all spam emails are considered to be phishing. For this reason, we only focused on the phishing email dataset.
3 Methodology
In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with the explainable AI. Initially, we collected a dataset and prepared it with data cleaning and balancing techniques. After preparing the data we have focused on optimizing and fine-tuning the model. For this phishing email detection purpose, we have selected the DistilBERT model and pre-trained with the prepared data. The DistilBERT model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts. At the end of the experiment, we employed two XAI techniques Explanations LIME and Transformer Interpret to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification. The overall methodology of the proposed method is depicted in Fig. 1.
Figure 1: The proposed overall methodology.
3.1 Data Collection and Processing
For this experiment, a publicly available dataset has been collected from kaggle. The dataset describes different kinds of email text bodies that can be used to identify phishing emails using in-depth text analysis. This dataset consists of “Safe Email” and “Phishing Email”. Fig. 2 shows the data overview. The dataset has 11,322 Safe Emails and 7,328 Phishing Emails which shows that it is an imbalanced dataset. Along with this, this dataset showed that null values exist in rows. As null values can’t be processed that’s why those rows are removed. By using the resampling technique minor class “Phishing Email” is oversampled. This resampling technique mainly generates copies of other instances and adds those copies to the dataset. It is essential to address uneven data in order to develop equitable and precise predictive models. Balancing data ensures algorithms work well across all categories by reducing biases toward the majority class. This prevents skewed outputs and makes it possible to create a more accurate and equitable model for underrepresented classes. After oversampling, the dataset has 11,322 safe emails and 11,322 phishing emails in equal proportion. This data balancing also helps to reduce the risk of oversampling. Fig. 3 illustrates the imbalance and balanced data before and after sampling.
Figure 2: The sample data overview
(a) Imbalanced data
(b) Balanced data
Figure 3: Imbalanced and balanced data.
3.2 Dataset splitting
For this experiment, we have split the dataset into 70-30% where 70% data is training data and 30% is testing data. In 70% training set total of 15,850 data exist and the testing set consists of 6,794 data. Initially, the training set was utilized for the train the model, and the testing test was used to find out the effectiveness of the model on unseen data identification and prediction is phishing or not.
3.3 Model Selection
In this experiment, we have employed the transformer-based BERT variant architecture DistilBERT which is a lightweight, fast, and economical model. This is developed for language processing tasks and also compresses the original BERT model and makes it faster. Knowledge distillation was utilized in the pre-training phase of this pre-trained model, which allowed it to be effectively reduced in size by 40% while maintaining 97% of its language understanding skills and delivering a 60% speed increase (Sanh et al., 2019). This small variant has twelve attention heads, 768 hidden layers, and six transformer layers (Adoma et al., 2020). Whereas DistilBERT contains about 66 million parameters, the original BERT model has 110M (million) parameters. By lowering the number of parameters, DistilBERT maintains excellent performance while achieving a smaller size and faster inference speed.
3.3.1 Model Optimization
To improve its handling of terms that are not in its lexicon and to capture meaningful sub-word representations, BERT uses a Word-Piece tokenizer, which splits words into smaller sub-words. The preprocessed phishing dataset is tokenized using the tokenizer of the DistilBERT transformer. In this tokenized portion, the maximum sequence length of input data has been used 512. Other crucial components of tokenization are padding and truncation. The major purpose of these procedures is to ensure that each tokenized input sequence has a constant length. This is important because it creates consistent embeddings and efficient processing in batches.
Pre-trained weights acquired throughout the pre-training phase are used to initialize the DistilBERT models. Selecting an optimal batch size is paramount in deep learning, as it significantly influences both training and testing accuracies, along with overall runtime (Lin, 2022). The careful selection of batch size determines the efficiency of the neural network training process, striking a balance between computational resources and model performance. The batch sizes for training and testing data are optimized at 32 and 64, respectively. Each epoch’s training data is assigned at random to ensure the model sees a diversity of unseen data samples and prevents overfitting. This shuffling method increases the model’s ability to generalize.
During training, an optimizer is essential in adjusting the parameters of a neural network. The main goal of this is to minimize the error or loss function of the model for better performance. For this research work, AdamW (Adam Weight Decay) (Loshchilov and Hutter, 2017) optimizer has been used to get better performance and avoid overfitting. This optimizer is a modification version of Adam which is widely adopted for training deep neural networks (Zhuang et al., 2022). By separating weight decay from the learning rate, AdamW enhances its implementation. It improves regularization and training dynamics by separating weight decay from gradient updates, in contrast to Adam. Learning rate is a crucial hyperparameter governing step size during training for optimizing models. The convergence and overall performance of the system are dependent on various aspects such as the model architecture, optimization techniques, and task domain. For efficient fine-tuning and fruitful model training, this parameter must be adjusted. Selecting a high learning rate can introduce instability, resulting in poor performance on unseen data, while a low learning rate may extend the training process, requiring more epochs and increasing computational costs. In this research, a learning rate, 2e-5 is used in the training process for the transformer-based DistilBERT model.
The overall discussion about model optimization indicates the parameter settings for the introduced best model for this specific experiment. The hyperparameter settings for the suggested model are shown in Table 1. We worked with every possible value of the parameters by using an iterative process to find the ideal value. Following that, these improved hyperparameter settings are used to train the suggested model.
Table 1: Parameter Settings for Phishing Email Detection Task.
Parameter Name Worked With Optimal Value
Training Batch size 8, 16, 32 32
Testing Batch size 8, 16, 32, 64 64
Optimizer AdamW AdamW
Learning rate 1e-5, 2e-5, 3e-5 2e-5
Number of epochs 3, 4, 5, 6, 7 6
3.3.2 Fine Tuning
A pre-trained model is fine-tuned using task-specific data and backpropagation to alter parameters on certain tasks and datasets. The primary objective of this is to improve the model’s performance across various tasks and datasets while minimizing the loss function. In this experiment, DistilBERTForSequenceClassification has been utilized to adopt the pre-trained model. DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command where train_loader is configured with shuffling and test_loader is configured without shuffling. After that, we fine-tuned this DistilBERT model by adjusting and adopting some necessary steps which are shown in Fig.  4.
Figure 4: The overall process of fine-tuning.
In this overall fine-tuning process, following the optimizer’s setup, the model is fine-tuned by feeding it the input parameters (input_ids, attention_mask, and label) during the forward pass so that predictions can be made. After the forward pass, gradients of the loss for the parameters (weights and biases) are calculated in the backward pass. These gradients show how changing the parameters would affect the loss. Lastly, an update is made to the model parameters. The optimizer is mainly responsible for changing the neural network’s parameters in accordance with the calculated gradients. It adjusts the weights and biases in an attempt to reduce the loss. The model gains the capacity to fine-tune its parameters throughout several epochs, which improves its predictive capability on the training set. The ultimate goal is to identify underlying patterns in the training set and use those patterns to generalize well to new data. Algorithm 1 shows the overall fine-tuning process in training.
Algorithm 1 Fine-Tuning the DistilBERT model
1:DistilBERT model, training data loader, optimizer, and number of epochs
2:Trained fine-tuned model
3:$\text{epochs}\Leftarrow n$
[…]
5:for epoch in range(epochsdo
6:     model.train()
7:     for batch in train_loader do
[…]
10:         $\text{labels}\Leftarrow\text{batch}[2]$
11:         $\text{gradients}\Leftarrow\text{0}$
[…]
13:         loss_calculation()
14:         backwardpass()
15:         update_parameters()
16:     end for
17:end for
3.3.3 Model Explainability
The ability of a model to explain its decisions is crucial since it provides insight into the process. Verifying the accuracy of judgments and reducing errors are made feasible by understanding how models work. Explainable AI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022). This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions. In this research work, we have worked with model explanability to solve the black box problem. To understand the model prediction two explainability models LIME and Transformers Interpret have been utilized.
Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box DistilBERT model. The purpose of the LIME method is to faithfully explain any classifier or regressor’s predictions (Ribeiro et al., 2016). For improved interpretability, LIME provides an understandable depiction of the decision-making process, providing light on the “why" behind the predictions or suggestions made by these models. In order to achieve model explainability, LIME first perturbs the email content by making modest adjustments, rearranging words, or adding new words. After that, the black-box transformer-based model is applied to the altered email content to generate predictions. From these impacted instances, LIME derives features. This local model should roughly represent the behaviour of the black-box model surrounding the chosen instance. Next, explanations for the black-box model’s prediction for the original data are produced using the local model. This explanation usually highlights the words or key elements most important during the model’s decision-making.
A Transformers Interpret tool was also utilized to explain the transformer-based model for this classification work. The class explainer process was carried out using this tool using a model and tokenizer. With the use of this class explanation, word attributions are computed. Positive attribution numbers signify a word’s positive contribution to the projected class, while negative numbers signify a word’s negative contribution. The terms that demonstrate a strong contribution to decision-making are also highlighted in this explanation. The overall process of model explainability is shown in Fig. 5. Algorithm 2 shows the explanation process of the model prediction.
Figure 5: The overall process of model explainability
Algorithm 2 Explainability Analysis
1:Trained model, tokenizer, and text
2:Explanation of the model prediction
3:class_names $\Leftarrow$ [X, Y, …N]
4:function Explain_Prediction(model, tokenizer, text)
5:     LIME_Explanation(model, text)
6:     Transformer_Interpret(model, tokenizer, text)
7:end function
8:function LIME_Explanation(model, text)
9:     explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow class\_names$)
10:     explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
11:     exp_class $\Leftarrow$ explanation.available_labels()
12:     exp $\Leftarrow$ explanation.as_list($\Leftarrow$ exp_class)
13:     explanation_visualize()
14:end function
15:function Transformer_Interpret(model, tokenizer, text)
16:     cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
17:     word_attributions $\Leftarrow$ cls_explainer(text)
18:     cls_explainer_visualize()
19:end function
4 Results
The fine-tuned transformer-based model was trained and tested with both unbalanced and balanced datasets. To assess the performance of the model different measurements such as confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated. Overall summarize the performance, a model needs to calculate the confusion matrix first. It is a kind of table that provides a detailed breakdown of the actual and predicted classification for specific task domain data. The confusion matrix of the DistilBERT model for both imbalanced and balanced datasets is shown in Table 2. True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix and as well as are needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022). The precision, recall, and F1-score of the DistilBERT model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
Table 2: Confusion matrix of DistilBERT for both imbalanced and balanced dataset
Imbalanced Balanced
Class Name Safe Email Phishing Email Safe Email Phishing Email
Safe Email 3235 116 3281 98
Phishing Email 24 2216 5 3410
Table 3: The precision, recall, and f1-score for imbalanced dataset.
Class Precision Recall F1-score Support
Safe Email 0.99 0.97 0.98 3351
Phishing Email 0.95 0.99 0.97 2240
Table 4: The precision, recall, and f1-score for balanced dataset.
Class Precision Recall F1-score Support
Safe Email 1.00 0.97 0.98 3379
Phishing Email 0.97 1.00 0.99 3415
Precision: Precision which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
Precision $=\frac{TP}{TP+FP}$ (1)
Recall: It calculates the ratio of true positives to all predicted positives to assess a model’s accuracy in identifying positive samples. It denotes the sensitivity of the model, capturing the actual positive rate in the class observations.
Recall $=\frac{TP}{TP+FN}$ (2)
F1-score: A balanced statistical metric that assesses a model’s efficaciousness in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall. It offers a thorough perspective that balances recall and precision, impacting the final score determined by the model’s projections of the positive rate and actual positive cases.
F1-score $=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
The frequency of valid predictions in a model is quantified by accuracy, which also represents the percentage of correctly predicted samples out of all the samples the model generates.
$\text{Accuracy}=\frac{\text{TN}+\text{TP}}{\text{TN}+\text{TP}+\text{FN}+\text{FP}}$ (4)
The accuracy of a model can differ based on the various parameters, data balanced or not. Initially in this experiment, an imbalanced dataset was utilized to assess the model’s performance in both training and testing. This dataset is highly imbalanced based on the safe email and phishing email data where the number of data of safe email is quite higher than phishing email. To train the model with 70% imbalanced data we have used 6 epochs with a batch size of 32. After 6 epochs the model showed 98.90% training accuracy for this imbalanced dataset with 0.0201 training loss. On the other hand, we get a testing accuracy 97.50 with testing loss 0.0900 with batch size 64. Fig.  6 shows a graph of accuracy and loss with 6 epochs for an imbalanced dataset.
Figure 6: Training vs. Testing Accuracy and Loss on Imbalanced Data
Moreover, we also conducted this experiment with a resampled and balanced dataset. After preparing the balanced dataset again used 70% of the data for training purposes. Using this portion of data, we worked on model optimization and fine-tuning and trained the model. This time with balanced data the fine-tuned model showed 99.07% training accuracy with 0.0218 training loss. After that, we tested this trained model with 30% unseen balanced data and got 98.48% testing accuracy with 0.0447 testing loss. This performance showed that the performance of the model with a balanced dataset is quite better than an imbalanced dataset. Fig. 7 shows a graph of accuracy and loss with 6 epochs for a balanced dataset. The performance comparison of the model on both balanced and imbalanced dataset shows in Table 5.
Figure 7: Training vs. Testing Accuracy and Loss on Imbalanced Data.
Table 5: The Performance comparison of fine-tuned DistilBERT model on Imbalanced and Balanced dataset.
Dataset Mode
Training Loss
Training Accuracy
Testing Loss
Testing Accuracy
Imbalanced
0.0201
98.90
0.0900
97.50
Balanced
0.0218
99.07
0.0447
98.48
Finally, we have compared our Phishing Email Detection study with some previous studies in the same field in order to assess its quality. This comparison study is summarized in Table 6, which evaluates our work from several angles. These include the models used for machine and deep learning, the datasets used, and whether or not the study offers human interpretation and decision explainability for the model.
Table 6: Explainability of Different Studies on Phishing Email Detection
Reference
Datasets
Models
Explainability Analysis
2023(Jamal et al., 2023)
Email Spam Detection Dataset, PhishingEmailData
DistilBERT, RoBERTA
No
2021(Alhogail and Alsabih, 2021)
Fraud Dataset
GCN
No
2019(Fang et al., 2019)
IWSPA-AP
THEMIS
No
2023(Thapa et al., 2023)
IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl
THEMIS, BERT
No
2023(Atawneh and Aljehani, 2023)
Phishing Email Dataset, Benign Email Dataset
BERT-LSTM
No
This Study Phishing Email Fine-tuned DistilBERT Yes
The comparison showed that our research work provides transparency in this phishing email detection task. However, some listed research work showed higher accuracy than our proposed model. This happened because they used different datasets in their experiment. In this study, we only focused on the phishing email dataset but other existing works were also done this work with the spam dataset. Furthermore, to the best of our knowledge, none of the papers listed in Table  6 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.
5 Explainability Analysis
LIME and Transformer Interpret both showed how the fine-tuned DistilBERT model arrives at its prediction on phishing email detection by approximating the model’s behaviour on every word of the email text. Lime and Transformer Interpret provided accurate results by answering the “Why” model predicting the selected text is a “Safe Email” or “Phishing Email”. Based on the explainability, those explainable techniques generate word attributions by providing positive and negative numbers on each word. Those numbers showed which word was more specifically reasonable for the predicted class. By using those numbers the explainable model visualizes the whole text by highlighting the words that scored positive and as well as negative on the prediction.
For this explainability analysis in LIME explainer, we initialized the number of features (words) 15 for the dimensionality reduction. If the number of words of a selected test is very large that time the computational time will increase as well as it increases the dimensionality which also can be a reason why the systems crashed. Limiting this number of features makes the explanation process computationally more feasible and provides a clearer and more concise picture of which words in the case of text are influential in the model’s prediction for a particular instance. Fig. 8 shows the model explainability using LIME and Table  7 and Table  8 show the positive and negative weights associated with words that help users understand phishing email detection.
(a) Explanation for the first text
(b) Explanation for the second text
Figure 8: LIME Explanation for two Phishing Email Texts
Table 7: The positive and negative weights using LIME for first text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 1.760212017947662 ‘neonate’ -0.1267345497528774
‘price’ 1.5916451492409904 ‘3107’ -0.013956063314091143
‘software’ 1.2586059555064724
‘setupmefree’ 0.9966181271429526
‘com’ 0.3984979685547945
‘http’ 0.1447738056917774
‘a’ 0.06584106490791478
Table 8: The positive and negative weights using LIME for second text
Positive Numbers Negative Numbers
Word Value Word Value
‘megapixel’ 0.45124215111245936 ‘a’ -0.09469945665856638
‘get’ 0.4369880334209157
‘free’ 0.4206519368628723
‘digital’ 0.3721259726068145
‘camera’ 0.2388125335048898
‘canon’ 0.19498844421129996
‘6’ 0.04953986530836751
‘3’ 0.024635744735869915
Along with LIME, we also examined the model interpretability with similar text by using Transformer Interpret. During this interpretation, this Transformer Interpret uses the tokenizer to split the large words into sub-words with the ‘##’ prefix. This happens when an unusual or non-vocabulary word is encountered by the model, which breaks it down into smaller sub-words to enable efficient processing. Fig. 9 shows Transformer Interpretation for a phishing email and Table  9 and Table  10 show the word importance, with positive and negative numbers associated with each word for two phishing emails.
(a) Explanation for first text
(b) Explanation for second text
Figure 9: Transformer Interpret Explanation for two Phishing Email Text
Table 9: The positive and negative numbers using Transformer Interpret for first Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘low’ 0.7698059423920272 ‘.’ -0.09537831853568463
‘price’ 0.4465021887746993 ‘setup’ -0.037898186884516454
‘software’ 0.10363795047092349 ‘?’ -0.04837875144898304
‘http’ 0.05008396373925551 ‘a’ -0.212114259929307
‘:’ 0.06724251662885174 ’=’ -0.13427331689723981
‘/’ 0.018869178109329452 ‘310’ -0.06162690120821821
‘neon’ 0.07499325538696087 ‘##7’ -0.013360917838605864
‘##ate’ 0.09833698869154901 ‘.’ -0.1269640647507237
‘me’ 0.09196861346171359
‘free’ 0.16949987141093556
‘com’ 0.20340796716052334
‘/’ 0.006729002810807768
Table 10: The positive and negative numbers using Transformer Interpret for second Phishing Email Text
Positive Numbers Negative Numbers
Word Value Word Value
‘get’ 0.49733595483460524 ‘a’ -0.2688505206057748
‘free’ 0.4645088664478948 ‘6’ -0.09532782298942542
‘mega’ 0.18191195439136695 ‘.’ -0.08192496249808812
‘3’ 0.014327378740528552 ‘##pi’ -0.11609630028309895
‘##x’ 0.5855798135703232 ‘!’ -0.18838964048006548
‘digital’ 0.021290989158880363
‘canon’ 0.13188078942699152
‘camera’ 0.05155301993306188
‘>’ 0.03909299350598931
In Fig. 9 the highlighted text with green color shows the word importance to classify the analyzed text as Phishing. On the other hand, the red portions showed those words or sub-words that worked against the classified class. As we worked with a binary classification task here “True Label" = 1 means "Phishing Email" and when “True Label" = 0 means “Safe Email”.
According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the two tested input texts, but they differ in how they define explanation. LIME mostly uses small textual modifications to evaluate the impact of the model’s prediction on the entire text, while Transformer Interpret employs an attention mechanism to examine each segment of the input sequence. Furthermore, in the LIME and Transformer Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email contents. The positive words are mainly responsible for the predicting class. More specifically, when those words are present in a text that time model predicts a class. A comprehensive difference between the two explainable AI techniques is shown in Fig. 10 and Fig. 11. The number of scoring percentages for those email texts is not the same for LIME and Transformer Interpret. This is only because the working procedures of those techniques are not the same.
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 10: LIME and Transformer Interpret Comparison on first text
(a) LIME Scoring
(b) Transformer Interpret Scoring
Figure 11: LIME and Transformer Interpret Comparison on second text
6 Discussion
The results of both an imbalanced and balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned DistilBERT model. The fine-tuned version of the DistilBERT model showed 98.90% training and 97.50% testing accuracy on an imbalanced dataset whereas this model showed 99.07% training and 98.48% testing accuracy on the balanced dataset. Although these results show that the model’s performance on a balanced dataset is quite better than an imbalanced dataset, the difference is not so high, approximately 1%. This demonstrated that the fine-tuned version performs well in an imbalanced dataset. Additionally, the performance shows that the optimized model performed better than other previous studies.
Besides, this research work also answered the “How was the model performed?" and “Why was it performed?" using XAI, motivated from recently published book by Sarker (Sarker, 2024). Two well-known Explainable and interpretable techniques are utilized for the model understanding. It is more difficult to understand how DistilBERT makes its predictions because, as we know, it is a complex model. These AI techniques aided in gaining an understanding of the model’s decision-making process. Both LIME and Transformer Interpret highlighted the areas of the text that are more responsible for the classification of the phishing of safe email. This highlights areas worked with words or sub-words, making the decision more understandable for humans, and users can also comprehend it easily. Mainly, those techniques calculated each word or sub-words numeric value or weight where some words indicate positive value and some negative. Words with calculated higher positive values are more important for classifying the predicted class, while those with more negative values do not support the predicted class. This experimental analysis showed the transparency of the model in this phishing email classification task.
7 Conclusion
This study has investigated the detection of phishing emails by utilizing a Large Language Model. This also has shown how effective it is to use state-of-the-art natural language processing methods to improve the precision and effectiveness of phishing email identification. The overall experiment was done with both a balanced and imbalanced dataset. We have obtained promising results in differentiating between phishing and valid emails by utilizing DistilBERT’s contextual language understanding technique. Mainly, to achieve this good performance the model is optimized by changing some important parameters such as batch size, optimization function, learning rate, etc, and also fine-tuning has been done. This fine-tuned version of DistilBERT obtained 98.48% testing accuracy in this phishing email classification task. Furthermore, LIME and Transformer Interpret techniques have been employed to see the explanation of the model and how it makes decisions. The decision-making process is now more visible and understandable for users because of the usage of explainable AI, which promotes comprehension and trust. The overall model explainability describes the robustness of the proposed model.
The field of Large Language Models has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers. New open-source models are being produced almost every day, demonstrating the ongoing development and innovation in this field. In the future, we plan to expand the scope of our research by experimenting with a wider variety of models and datasets. Furthermore, we intend to investigate the use of different variants of BERT-based models for the identification of different language phishing emails. Moreover, deep analysis of the model will be conducted to ensure its robustness by enhancing the explainable AI techniques.
[…]
Data Availability Statement
The dataset that is used in this study is available at Kaggle (Online Link):  https://www.kaggle.com/datasets/subhajournal/phishingemails.

### v2

@@
-Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation
+Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Conceptualization, Writing - Original draft preparation
 Abstract
 Phishing email is a serious cyber threat that tries to deceive users by sending false emails with the intention of stealing confidential information or causing financial harm.
-Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make detection and prevention of phishing more challenging.
+Attackers, often posing as trustworthy entities, exploit technological advancements and sophistication to make the detection and prevention of phishing more challenging.
 Despite extensive academic research, phishing detection remains an ongoing and formidable challenge in the cybersecurity landscape.
 Large Language Models (LLMs) and Masked Language Models (MLMs) possess immense potential to offer innovative solutions to address long-standing challenges.
-In this research paper, we present an optimized, fine-tuned transformer-based DistilBERT model designed for the detection of phishing emails.
-In the detection process, we work with a phishing email dataset and utilize the preprocessing techniques to clean and solve the imbalance class issues.
-Through our experiments, we found that our model effectively achieves high accuracy, demonstrating its capability to perform well.
-Finally, we demonstrate our fine-tuned model using Explainable-AI (XAI) techniques such as Local Interpretable Model-Agnostic Explanations (LIME) and Transformer Interpret to explain how our model makes predictions in the context of text classification for phishing emails.
+In this research paper, we present a fine-tuned transformer-based language model, RoBERTa (Robustly Optimized BERT Pretraining Approach), for detecting phishing emails.
+In the detection process, we employ a real-world phishing email dataset and apply the preprocessing techniques to clean and address the class imbalance issues, thereby enhancing model performance.
+Our experimental results demonstrate that the fine-tuned model achieves high accuracy, confirming its robustness and effectiveness in the phishing email classification task.
+To ensure model transparency and trust in the model’s decision-making, we propose a hybrid explanation approach, LITA (LIME-Transformer Attribution), which integrates the potential of Local Interpretable Model-Agnostic Explanations (LIME) and Transformers Interpret methods.
+This hybrid method provides more user-friendly insights and solves the disagreement issues between the two approaches.
+Moreover, we also present how the model generates its predictions by presenting positive and negative contribution scores using LIME, Transformers Interpret, and LITA.
 keywords:
@@
 The rapid advancement of internet technology has significantly changed the way people communicate online while also raising more security issues (Salloum et al., 2021).
-One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation.
+One of the prevalent security concerns, the cyberattack termed “email phishing” poses grave risks such as identity or login credentials theft, financial repercussions, and harm to an organization’s reputation (Almomani et al., 2013).
 In this current decade, email phishing has changed throughout time to become a more effective method for carrying out fraudulent actions.
@@
 To evade falling victim to this threat, organizations need to adeptly detect and prevent these attacks at scale by adopting effective phishing detection techniques.
-Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023).
+Leveraging Artificial Intelligence (AI) techniques like Machine Learning, Deep Learning, and Hybrid Learning, which are already utilized for phishing attack detection, can significantly bolster these efforts (Basit et al., 2021) (Sarker, 2023).
 Moreover, because of the rapid advancements in technology, new methods based on AI are being developed daily in the cybersecurity field.
-Transformer-based models become more popular in the context of text classification.
+Transformer-based models have become more popular in the context of text classification.
 The transformer is an innovative neural design that uses the attention mechanism to encode incoming data into powerful characteristics (Han et al., 2021).
 Transformer-based models have greatly improved the processing, comprehension, and interpretation of text data inputs, which has had a revolutionary effect on the development of phishing classification techniques (Jamal et al., 2023).
-Large Language Models like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023).
+LLMs like OpenAI’s ChatGPT (Generative Pre-trained Transformer) and Google’s Bard have transformed the landscape of natural language creation and comprehension, capturing considerable attention from the general public (Zhao et al., 2023; Yao et al., 2023).
 The increasing complexity of LLMs is radically changing artificial intelligence and has the potential to transform algorithm development and application by understanding human language through extensive text data patterns and correlations.
-Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as Large Language Models because of their ability to comprehend and produce text that is similar to that of a human.
-BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to grasp contextual word embeddings (Koroteev, 2021), become increasingly efficient and valuable in text classification and generation, including tasks like spam and phishing detection, sentiment analysis (Singh et al., 2021), and more other fields.
-BERT is designed to take into account both left and right contexts at every level to pre-train complex bidirectional representations from unlabeled text (Devlin et al., 2018).
+Many transformer-based models used in natural language processing (NLP), including well-known ones like OpenAI’s GPT and Google’s BERT (Bidirectional Encoder Representations from Transformers), are categorized as language models due to their large number of parameters and are designed to process and generate human language.
+The BERT model, a type of transformer-based neural network architecture pre-trained on extensive text corpora to capture contextual word embeddings (Koroteev, 2021), has become increasingly efficient and valuable in text classification and generation tasks, including spam and phishing detection, sentiment analysis (Singh et al., 2021), and various other fields.
+BERT utilizes a bidirectional representation approach that considers both left and right context simultaneously at all layers, allowing it to learn rich language representations from large amounts of unlabeled text (Devlin et al., 2018).
 With simply one extra output layer, this method allows the pre-trained BERT model to be refined, leading to the development of state-of-the-art models for a variety of applications.
@@
 BERT makes use of a Transformer architecture technique called “self-attention”, which allows it to recognize inter-word relationships in text sequences and assess the relative importance of individual words within a sentence in order to understand the context.
-Based on the transformer layers number and parameters BERT model has different variants such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
+Based on the transformer layers number and parameters BERT model has different variants, such as: BERT-base (Khadhraoui et al., 2022), ALBERT (Lan et al., 2019), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2019), etc.
 Despite BERT’s versatility across NLP tasks, comprehending its internal mechanisms often poses challenges, resembling a “black box” due to its complexity.
@@
 XAI techniques emerge as a vital solution, offering significant perceptions of the model’s reasoning and illuminating the decision-making process (Hoffman et al., 2018; Anan et al., 2023).
-Recently, Sarker (Sarker, 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested to explore more with experiments.
-XAI techniques can illuminate the words or features deemed most influential by the BERT model in specific text classification tasks of any nature.
-In this research, our main goal is to show transparency in phishing email detection with explainable AI.
-The major contribution of the work is as follows:
-  • We present a transformer-based Phishing email detection model, a fine-tuned version of DistilBERT.
-  • Two different XAI techniques LIME and Transformer Interpret are utilized to explore the interpretability of the proposed fine-tuned model and visualize the scoring percentage of those techniques on prediction.
-  • We demonstrate our model performance on both imbalanced and balanced datasets.
+Recently, Sarker (Sarker, 2024; Sarker et al., 2024) has also highlighted the significance of explainability analysis in various perspectives in the context of cybersecurity, in which we are interested in exploring more with experiments.
+Explainable methods can highlight the words, sub-words, or features that the language model considers most influential in various text classification tasks.
+In our previous work (Uddin et al., 2025), we implemented two explainable methods and provided explanations with positive and negative coefficients of words.
+In this research, our main goal is to focus transparency in phishing email detection with explainable AI and to propose a hybrid method that considers the limitations of existing approaches.
+The major contributions of the work are as follows:
+  • We build a robust phishing email detection model by fine-tuning a pre-trained RoBERTa transformer-based language model.
+  • We present a hybrid explanation approach, LITA, to provide a more user-friendly explanation for transformer-based phishing email detection.
+In addition, to produce alternate explanations, we also use LIME and Transformers Interpret, which enables us to compare results and draw attention to the limitations of current XAI methods.
+  • We calculate each word’s positive and negative contribution score in a phishing email to show how it affects the phishing classification result.
+  • We evaluate our method on a real-world dataset of phishing emails, showcasing the efficacy of the model and the clarity of the explanations it generates.
 The remaining sections of the paper are structured as follows: Section 2 delves into related research works.
@@
 Cyber-attacks of all kinds are becoming more and more frequent, and they can destroy or steal imported data.
-In addition, in this cybersecurity field, there are an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks.
+In addition, in this cybersecurity field, there is an enormous amount of new and creative techniques being developed daily for identifying and preventing cyberattacks.
 For the improvement of phishing email detection and prevention systems, many researchers have developed a wide range of novel strategies to counteract phishing emails, utilizing state-of-the-art methods from deep learning, machine learning, and hybrid models.
 Phishing is still a serious cybersecurity concern and one of the main ways hackers breach networks.
+2.1 Classical Machine Learning Methods
 Detecting phishing attempts early on is critical for businesses nowadays.
@@
 Identifying the behaviour-based features in phishing emails was the main objective of this research because an attacker cannot disguise these types of features.
-They examined the overall process with a phishing dataset that consists of 6923 emails where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset and achieved 94% accuracy with this hybrid feature selection technique.
+They examined the overall process with a phishing dataset that consists of 6923 emails, where phishing emails were collected from the Nazario dataset and legitimate emails from the SpamAssassin dataset, and achieved 94% accuracy with this hybrid feature selection technique.
 With an emphasis on the importance of top-ranking features, the study (Zamir et al., 2020) examines feature selection methods (GR, IG, Relief-F, RFE) on a dataset of 11,055 phishing sites with 32 attributes.
@@
 A variety of supervised learning methods with differing feature sets are compared, such as RF, SVM, bagging, kNN, Neural Network (NN), and J48.
-The proposed method using RF showed 97.3% and using stacking achieved 97.4% accuracy.
-In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts.Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP).
-A publicly available balanced phishing dataset has been utilized where the authors used the email body’s text to extract features that improved the proposed model’s accuracy.
+The proposed method using RF showed 97.3%, and using stacking achieved 97.4% accuracy.
+Almomani et al.
+(2013) presented the Phishing Dynamic Evolving Neural Fuzzy Framework (PDENF), which detects and predicts unknown "zero-day" phishing emails with reduced false positive and false negative rates.
+It leverages hybrid and evolving connectionist systems for rapid, lifelong learning, featuring a minimal memory footprint and a simplified rule base for email classification.
+Salhi et al.
+(2021) demonstrated a novel email clustering approach that uses information gain techniques to evaluate email attributes, enabling effective differentiation between benign and harmful emails.
+2.2 Deep Learning Methods
+In cybersecurity, deep learning has great potential, particularly for detecting network intrusions, classifying malware, and recognizing phishing attempts.
+Alhogail and Alsabih (2021) proposed a deep learning-based phishing email detection model based on graph convolutional network (GCN) and natural language processing (NLP).
+A publicly available balanced phishing dataset has been utilized, where the authors extracted features from the email body’s text to improve the proposed model’s accuracy.
 The model showed 98.2% accuracy with a low false-positive rate of 0.015%.
@@
 Data preprocessing (email cleaning, tokenization, and stop-word elimination) techniques were employed to derive the important features from the Enron email dataset.
-The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameter of the proposed model and achieved a high accuracy of 99.72%.Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam email detection.
-They developed a phish responder which is a Python-based command line solution and used deep learning and NLP for this classification task.
+The Gated Recurrent Unit (GRU) model was used with a CS algorithm for the fine-tuning of the parameters of the proposed model and achieved a high accuracy of 99.72%.
+Dewis and Viana (2022) employed Long short-term memory (LSTM) model with NLP to detect both phishing and spam emails.
+They developed a phish responder, which is a Python-based command-line solution, and used deep learning and NLP for this classification task.
 The proposed LSTM model achieved 99% accuracy for the text-based dataset.
 Fang et al.
-(2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural networks (RCNN) model which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM).
-99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced where the positive rate (FPR) is 0.043%.
-Other deep learning algorithms, such the reinforcement learning, Recurrent Neural Network (RNN), and Deep Belief Network (DBN), are also in use (Zhang and Li, 2017; Bahnsen et al., 2017; Smadi et al., 2018).
+(2019) introduced THEMIS, a phishing email detection model built on an improved recurrent convolutional neural network (RCNN) model, which was improved by using the Bidirectional Long Short-Term Memory (Bi-LSTM).
+99.84% accuracy is demonstrated by the suggested THEMIS model on datasets that are not balanced, where the positive rate (FPR) is 0.043%.
+In addition, researchers investigated deep neural architectures such as Recurrent Neural Networks (RNNs) and Deep Belief Networks (DBNs) (Zhang and Li, 2017; Bahnsen et al., 2017) in the context of phishing detection, as well as a Neural Network (NN) based on Reinforcement Learning (RL), called DENNuRL, for adaptive decision-making (Smadi et al., 2018).
+2.3 Transformer-based Methods
 Moreover, transformer-based models are also making a good impact in the cybersecurity field.Thapa et al.
@@
 The proposed model using BERT and LSTM, reaching an impressive accuracy of 99.61% on publicly available 4 different datasets, outperforming other methods.
-Nowadays, other variants of the BERT model such as DistilBERT, RoBERTa, and TinyBER (Jiao et al., 2019) are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023).
-The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants such as the fine-tuned DistilBERT and RoBERTA model and showed good accuracy in this research scope.
+Nowadays, other variants of the BERT model, such as DistilBERT, RoBERTa, and TinyBERT (Jiao et al., 2019), are performing well in the phishing and spam detection research sector (Jamal et al., 2023; Lee et al., 2020; Songailaitė et al., 2023).
+The authors (Jamal et al., 2023) presented an Improved Phishing and Spam Detection Model (IPSDM) which consists of two BERT variants, such as the fine-tuned DistilBERT and RoBERTA model,s and showed good accuracy in this research scope.
 Lee et al.
 Lee et al.
-(2020) proposed Context-Aware Tiny Bert (CatBERT) which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos.
-The detection rate of this proposed model was 87% which was far better than other compared models.
+(2020) proposed Context-Aware Tiny Bert (CatBERT), which exhibits faster processing and robustness to adversarial attacks than other transformer models, especially those that involve intentional keyword substitutions with synonyms or typos.
+The detection rate of this proposed model was 87%, which was far better than other compared models.
 Furthermore, researchers also used fine-tuned BERT-based models for phishing URL detection (Wang et al., 2023; Maneriker et al., 2021).
-Overall, this literature survey shows the existing works although showing good performance by using different datasets and models but most of the research didn’t work model’s interpretability or explainability.
-“How does a model work?” or “Which things or materials contribute to detecting phishing?” those types of questions can arise in their working procedure.
+Overall, this literature survey shows that although the existing studies demonstrate good performance using different datasets and models, most of them do not address model interpretability or explainability.
+Questions such as “How does the model work?” or “Which features contribute to phishing detection?” can arise in their methodologies.
 In this work, to answer those types of questions and solve the existing drawbacks in the literature, we have worked with explainable AI using not only one technique but also used two different techniques with different methodologies.
-In the literature, although some researchers did full experiments for phishing email detection but they utilized spam datasets as phishing datasets.
+Although several studies have conducted comprehensive experiments on phishing email detection, many of them relied on spam datasets rather than datasets specifically designed for phishing, which may limit the validity of their findings.
 In cybersecurity, we know that two different types of emails exist: spam and phishing, and not all spam emails are considered to be phishing.
@@
 3 Methodology
-In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with the explainable AI.
+In this paper, we have tried to explore the Transformer-based self-attention mechanism BERT model with explainable AI.
 Initially, we collected a dataset and prepared it with data cleaning and balancing techniques.
-After preparing the data we have focused on optimizing and fine-tuning the model.
-For this phishing email detection purpose, we have selected the DistilBERT model and pre-trained with the prepared data.
-The DistilBERT model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts.
-At the end of the experiment, we employed two XAI techniques Explanations LIME and Transformer Interpret to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification.
-The overall methodology of the proposed method is depicted in Fig.
+After preparing the data, we have focused on optimizing and fine-tuning the model.
+For this phishing email detection purpose, we have selected the RoBERTa model based on our previous experimental analysis Uddin et al.
+(2025) and pre-trained it with the prepared data.
+The RoBERTa model demonstrates good performance in both balanced and imbalanced scenarios of datasets when identifying phishing emails from texts.
+At the end of the experiment, we employed XAI techniques to explore our fine-tuned model and described how it is working and performing with good accuracy in this text classification.
+The methodology of the proposed work is depicted in Fig.
 1.
-Figure 1: The proposed overall methodology.
-3.1 Data Collection and Processing
+Figure 1: An overview of our experimental methodology.
+3.1 Data Characteristics and Processing
 For this experiment, a publicly available dataset has been collected from Kaggle.
 The dataset describes different kinds of email text bodies that can be used to identify phishing emails using in-depth text analysis.
-This dataset consists of “Safe Email” and “Phishing Email”.
+This dataset consists of two different types of email: Legitimate or Legitimate Email and Phishing Email.
 Fig.
-2 shows the data overview.
-The dataset has 11,322 Safe Emails and 7,328 Phishing Emails which shows that it is an imbalanced dataset.
-Along with this, this dataset showed that null values exist in rows.
-As null values can’t be processed that’s why those rows are removed.
-By using the resampling technique minor class “Phishing Email” is oversampled.
-This resampling technique mainly generates copies of other instances and adds those copies to the dataset.
-It is essential to address uneven data in order to develop equitable and precise predictive models.
+2 shows the phishing email samples.
+The dataset comprises 11,322 legitimate Emails and 7,328 phishing Emails, indicating that it is imbalanced.
+Additionally, this dataset revealed the presence of null values in certain rows.
+As null values can’t be processed, that’s why those rows are removed.
+We initially focused on data cleaning before addressing the data imbalance issue.
+This process involved converting the input text to lowercase, removing HTML tags that may be present in the text, replacing multiple whitespace characters with a single space, removing carriage return characters, and trimming leading and trailing whitespace.
+These steps ensured that the text was in a clear and consistent format for further processing.
+After cleaning the data, we utilized a data augmentation technique called synonym replacement to address the imbalance problem.
+This technique replaces random words from the original text with their synonyms, preserving the original meaning of the sentence while maintaining its similarity to the original text.
+In this context, we randomly selected words that are not common stopwords and replaced them with their randomly chosen synonyms.
+By using the augmentation technique, the minor class “Phishing Email” is oversampled.
+It is essential to address uneven data to develop equitable and precise predictive models.
 Balancing data ensures algorithms work well across all categories by reducing biases toward the majority class.
 This prevents skewed outputs and makes it possible to create a more accurate and equitable model for underrepresented classes.
-After oversampling, the dataset has 11,322 safe emails and 11,322 phishing emails in equal proportion.
+After oversampling, the dataset has 11,322 legitimate emails and 11,322 phishing emails in equal proportion.
 This data balancing also helps to reduce the risk of oversampling.
 Fig.
-3 illustrates the imbalance and balanced data before and after sampling.
-Figure 2: The sample data overview
+3 illustrates the imbalanced and balanced data before and after oversampling.
+Figure 2: The sample data of Phishing Emails
 (a) Imbalanced data
@@
 3.2 Dataset splitting
-For this experiment, we have split the dataset into 70-30% where 70% data is training data and 30% is testing data.
-In 70% training set total of 15,850 data exist and the testing set consists of 6,794 data.
-Initially, the training set was utilized for the train the model, and the testing test was used to find out the effectiveness of the model on unseen data identification and prediction is phishing or not.
+For this experiment, we split the dataset into 80-20%, where 80% of the data is used for training and 20% for testing.
+The 80% training data is then further split into 80-20%, resulting in an overall split of 64% training set, 16% validation set, and 20% testing set.
+The 64% training set contains a total of 14,492 data, the validation set consists of 3,623 data, and the testing set comprises 4,529 data.
+Initially, the training set was used to train the model, the validation set was used to validate the model to ensure its generalizability, and the testing set was used to evaluate the model’s effectiveness in identifying and predicting phishing on unseen data.
 3.3 Model Selection
-In this experiment, we have employed the transformer-based BERT variant architecture DistilBERT which is a lightweight, fast, and economical model.
-This is developed for language processing tasks and also compresses the original BERT model and makes it faster.
-Knowledge distillation was utilized in the pre-training phase of this pre-trained model, which allowed it to be effectively reduced in size by 40% while maintaining 97% of its language understanding skills and delivering a 60% speed increase (Sanh et al., 2019).
-This small variant has twelve attention heads, 768 hidden layers, and six transformer layers (Adoma et al., 2020).
-Whereas DistilBERT contains about 66 million parameters, the original BERT model has 110M (million) parameters.
-By lowering the number of parameters, DistilBERT maintains excellent performance while achieving a smaller size and faster inference speed.
+In this experiment, we have employed RoBERTa, a transformer-based robustly optimized variant of BERT, which retains BERT’s architecture and has refined the training process through key modifications.
+RoBERTa is trained on a much larger text dataset (160 GB of text) using 500,000 training steps with a large batch size of 8,000 per unit(Liu et al., 2019).
+The training process benefited greatly from dynamic masking, which updates the masking scheme during training instead of using static masking like BERT did during data preprocessing.
+This allows the model to learn a much more diverse set of training examples (from the same text) and also improves robustness.
+Moreover, the model removes the Next Sentence Prediction (NSP) task that BERT used, only uses MLM, finding it unnecessary and potentially less helpful to performance.
+Additionally, it uses a more efficient tokenization method, Byte-Level Byte Pair Encoding (BPE), similar to GPT-2, which helps handle rare words better.
+The RoBERTa-base contains about 125 million parameters.
+This variant has 12 attention heads, a 768-dimensional hidden state, and 12 transformer layers.
+The architecture of the RoBERTa model is illustrated in Fig.
+4.
+Figure 4: The architecture of the RoBERTa model
 3.3.1 Model Optimization
-To improve its handling of terms that are not in its lexicon and to capture meaningful sub-word representations, BERT uses a Word-Piece tokenizer, which splits words into smaller sub-words.
-The preprocessed phishing dataset is tokenized using the tokenizer of the DistilBERT transformer.
-In this tokenized portion, the maximum sequence length of input data has been used 512.
-Other crucial components of tokenization are padding and truncation.
-The major purpose of these procedures is to ensure that each tokenized input sequence has a constant length.
-This is important because it creates consistent embeddings and efficient processing in batches.
-Pre-trained weights acquired throughout the pre-training phase are used to initialize the DistilBERT models.
+RoBERTa employs a BPE tokenizer that is capable of splitting words into smaller sub-word units to better manage out-of-vocabulary terms and to allow for more meaningful sub-word representations.
+The preprocessed phishing dataset is tokenized using RoBERTaTokenizer, the default tokenizer for the RoBERTa transformer model.
+The maximum sequence length of the input data is configured to be a maximum of 512 tokens during the tokenization process, as this is the standard length for most transformer-based architectures, including RoBERTa.
+Moreover, as we are working with an email dataset, where the email texts are often long, setting the token length to less than 512 could result in a potential loss of important information.
+Other important steps in the tokenization process include padding and truncation: padding fills in sequences that are shorter than the maximum length with special token padding tokens, and truncation removes tokens from the end of sequences that are longer than the maximum length.
+These two tasks are important as they ensure input dimensions are standardized across the algorithm, so that embeddings are consistent and the batches of data can be processed efficiently during training and inference time.
+Pre-trained weights acquired throughout the pre-training phase are used to initialize the RoBERTa model.
 Selecting an optimal batch size is paramount in deep learning, as it significantly influences both training and testing accuracies, along with overall runtime (Lin, 2022).
@@
 For this research work, AdamW (Adam Weight Decay) (Loshchilov and Hutter, 2017) optimizer has been used to get better performance and avoid overfitting.
-This optimizer is a modification version of Adam which is widely adopted for training deep neural networks (Zhuang et al., 2022).
+This optimizer is a modified version of Adam, which is widely adopted for training deep neural networks (Zhuang et al., 2022).
 By separating weight decay from the learning rate, AdamW enhances its implementation.
 It improves regularization and training dynamics by separating weight decay from gradient updates, in contrast to Adam.
-Learning rate is a crucial hyperparameter governing step size during training for optimizing models.
+Learning rate is a crucial hyperparameter governing the step size during training for optimizing models.
 The convergence and overall performance of the system are dependent on various aspects such as the model architecture, optimization techniques, and task domain.
@@
 Selecting a high learning rate can introduce instability, resulting in poor performance on unseen data, while a low learning rate may extend the training process, requiring more epochs and increasing computational costs.
-In this research, a learning rate, 2e-5 is used in the training process for the transformer-based DistilBERT model.
+In this research, a learning rate, 2e-5 is used in the training process for the transformer-based RoBERTa model.
 The overall discussion about model optimization indicates the parameter settings for the introduced best model for this specific experiment.
@@
 Learning rate 1e-5, 2e-5, 3e-5 2e-5
-Number of epochs 3, 4, 5, 6, 7 6
+Number of epochs 3-10 5
 3.3.2 Fine Tuning
@@
 The primary objective of this is to improve the model’s performance across various tasks and datasets while minimizing the loss function.
-In this experiment, DistilBERTForSequenceClassification has been utilized to adopt the pre-trained model.
-DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command where train_loader is configured with shuffling and test_loader is configured without shuffling.
-After that, we fine-tuned this DistilBERT model by adjusting and adopting some necessary steps which are shown in Fig.
-4.
-Figure 4: The overall process of fine-tuning.
+In this experiment, RobertaForSequenceClassification has been utilized to adopt the pre-trained model.
+DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command, where train_loader is configured with shuffling and test_loader is configured without shuffling.
+After that, we fine-tuned this RoBERTa model during the training process by adjusting and adopting some necessary steps: optimizer setup, forward pass, loss calculation, backward pass (backpropagation), and updating model parameters, which are shown in Fig.
+5.
+Figure 5: The process of fine-tuning.
 In this overall fine-tuning process, following the optimizer’s setup, the model is fine-tuned by feeding it the input parameters (input_ids, attention_mask, and label) during the forward pass so that predictions can be made.
@@
 The ultimate goal is to identify underlying patterns in the training set and use those patterns to generalize well to new data.
-Algorithm 1 shows the overall fine-tuning process in training.
-Algorithm 1 Fine-Tuning the DistilBERT model
-1:DistilBERT model, training data loader, optimizer, and number of epochs
-2:Trained fine-tuned model
-3:$\text{epochs}\Leftarrow n$
-[…]
-5:for epoch in range(epochsdo
-6:     model.train()
-7:     for batch in train_loader do
-[…]
-[…]
-10:         $\text{labels}\Leftarrow\text{batch}[2]$
-11:         $\text{gradients}\Leftarrow\text{0}$
-[…]
-13:         loss_calculation()
-14:         backwardpass()
-15:         update_parameters()
-16:     end for
-17:end for
 3.3.3 Model Explainability
@@
 Verifying the accuracy of judgments and reducing errors are made feasible by understanding how models work.
-Explainable AI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022).
-This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions.
-In this research work, we have worked with model explanability to solve the black box problem.
-To understand the model prediction two explainability models LIME and Transformers Interpret have been utilized.
-Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box DistilBERT model.
+XAI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022).
+This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions Sarker (2024).
+In this research work, we have worked with model explainability to solve the black box problem.
+To understand the model prediction, two popular explainability models, LIME and Transformers Interpret, have been utilized.
+The overall process of model explainability is shown in Fig.
+6.
+Figure 6: The overall process of model explainability
+Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box RoBERTa model.
 The purpose of the LIME method is to faithfully explain any classifier or regressor’s predictions (Ribeiro et al., 2016).
 For improved interpretability, LIME provides an understandable depiction of the decision-making process, providing light on the “why" behind the predictions or suggestions made by these models.
-In order to achieve model explainability, LIME first perturbs the email content by making modest adjustments, rearranging words, or adding new words.
+To achieve model explainability, LIME first perturbs the email content by making modest adjustments, such as rearranging words or adding new ones.
 After that, the black-box transformer-based model is applied to the altered email content to generate predictions.
@@
 Next, explanations for the black-box model’s prediction for the original data are produced using the local model.
-This explanation usually highlights the words or key elements most important during the model’s decision-making.
+This explanation usually highlights the words or key elements most important during the model’s decision-making process.
 A Transformers Interpret tool was also utilized to explain the transformer-based model for this classification work.
-The class explainer process was carried out using this tool using a model and tokenizer.
+The class explainer process was carried out using this tool, using a model and tokenizer.
 With the use of this class explanation, word attributions are computed.
@@
 The terms that demonstrate a strong contribution to decision-making are also highlighted in this explanation.
-The overall process of model explainability is shown in Fig.
-5.
-Algorithm 2 shows the explanation process of the model prediction.
-Figure 5: The overall process of model explainability
-Algorithm 2 Explainability Analysis
-1:Trained model, tokenizer, and text
+However, while LIME and Transformers Interpret both provide a suitable understanding of a model’s decision-making processes, they differ fundamentally in their methods for explanation.
+LIME operates on word-level perturbations, whereas Transformers Interpret uses gradient-based attribution on subword tokens derived from the tokenizer.
+As a result, the outputs of these techniques, such as the number of highlighted tokens, the magnitude of scores, and their interpretive granularity, often vary, potentially leading to confusion for users.
+To solve this limitation and enhance the clarity and consistency of the explanations, we present a hybrid explanation approach called LITA.
+This method integrates the attribution scores from both LIME and Transformers Interpret into a unified explanation by computing a weighted combination of their normalized outputs.
+Algorithm  1 shows the explanation process of the model prediction using a hybrid approach to generate comprehensive explanatory scores.
+Here, $\alpha$ is a tunable parameter in the range $[0,1]$, controlling the influence of each technique.
+Both LIME (norm_lime_scores) and Transformers (norm_transformer_scores) attribution scores are first normalized to a common range [1,1][-1,1] to ensure consistency and comparability.
+As we are not sure which method is more accurate so we set $\alpha=0.5$ in this experiment, which assigns equal weight to both methods.
+In addition to reducing the overall differences between both attribution techniques, offering an indication of total feature importance with a hybrid attribution technique offers an additional layer of interpretability and reliability of the decision-making process.
+Algorithm 1 Hybrid Explainability Method
+1:Trained model, tokenizer, and input text
 2:Explanation of the model prediction
-3:class_names $\Leftarrow$ [X, Y, …N]
+3:class_names $\Leftarrow$ [X, Y, …, N]
 4:function Explain_Prediction(model, tokenizer, text)
-5:     LIME_Explanation(model, text)
-6:     Transformer_Interpret(model, tokenizer, text)
-7:end function
-8:function LIME_Explanation(model, text)
-9:     explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow class\_names$)
-10:     explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
-11:     exp_class $\Leftarrow$ explanation.available_labels()
-12:     exp $\Leftarrow$ explanation.as_list($\Leftarrow$ exp_class)
-13:     explanation_visualize()
-14:end function
-15:function Transformer_Interpret(model, tokenizer, text)
-16:     cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
-17:     word_attributions $\Leftarrow$ cls_explainer(text)
-18:     cls_explainer_visualize()
-19:end function
+5:  lime_scores $\Leftarrow$ LIME_Explanation(model, text)
+6:  transformer_scores $\Leftarrow$ Transformer_Interpret(model, tokenizer, text)
+7:  normalized_lime $\Leftarrow$ lime_scores scaled to [-1, 1]
+8:  normalized_transformer $\Leftarrow$ transformer_scores scaled to [-1, 1]
+9:  hybrid_scores $\Leftarrow$ Hybrid_Explanation(normalized_lime, normalized_transformer, $\alpha$)
+10:  return lime_scores, transformer_scores, hybrid_scores
+11:end function
+12:function LIME_Explanation(model, text)
+13:  explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow$ class_names)
+14:  explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
+15:  exp_class $\Leftarrow$ explanation.available_labels()
+16:  lime_scores $\Leftarrow$ explanation.as_list(label $\Leftarrow$ exp_class)
+17:  return lime_scores
+18:end function
+19:function Transformer_Interpret(model, tokenizer, text)
+20:  cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
+21:  transformer_scores $\Leftarrow$ cls_explainer(text)
+22:  return transformer_scores
+23:end function
+24:function Hybrid_Explanation(norm_lime_scores, norm_transformer_scores, $\alpha$)
+25:  for all word in norm_lime_scores $\cap$ norm_transformer_scores do
+26:   hybrid[word] $\Leftarrow\alpha\cdot$ norm_lime_scores[word] $(1-\alpha)\cdot$ norm_transformer_scores[word]
+27:  end for
+28:  return hybrid
+29:end function
 4 Results
-The fine-tuned transformer-based model was trained and tested with both unbalanced and balanced datasets.
-To assess the performance of the model different measurements such as confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated.
+A model’s generalization power can be assessed using various performance metrics, including precision, recall, F1-score, and accuracy, across different conditions (Chen et al., 2024).
+The fine-tuned transformer-based model was trained and tested with unbalanced and balanced datasets.
+To assess the performance of the model, different measurements such as the confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated.
 Overall summarize the performance, a model needs to calculate the confusion matrix first.
-It is a kind of table that provides a detailed breakdown of the actual and predicted classification for specific task domain data.
-The confusion matrix of the DistilBERT model for both imbalanced and balanced datasets is shown in Table 2.
-True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix and as well as are needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022).
-The precision, recall, and F1-score of the DistilBERT model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
-Table 2: Confusion matrix of DistilBERT for both imbalanced and balanced dataset
+This kind of table provides a detailed breakdown of the actual and predicted classification for specific task domain data.
+The RoBERTa model’s confusion matrix for imbalanced and balanced datasets is shown in Table 2.
+True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix, and as well as are also needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022).
+The precision, recall, and F1-score of the RoBERTa model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
+Table 2: Confusion matrix of RoBERTa for both imbalanced and balanced datasets
 Imbalanced Balanced
-Class Name Safe Email Phishing Email Safe Email Phishing Email
-Safe Email 3235 116 3281 98
-Phishing Email 24 2216 5 3410
-Table 3: The precision, recall, and f1-score for imbalanced dataset.
+Class Name Legitimate Email Phishing Email Legitimate Email Phishing Email
+Legitimate Email 2150 59 2181 60
+Phishing Email 17 1501 10 2278
+Table 3: The precision, recall, and f1-score for RoBERTa model on imbalanced dataset.
 Class Precision Recall F1-score Support
-Safe Email 0.99 0.97 0.98 3351
-Phishing Email 0.95 0.99 0.97 2240
-Table 4: The precision, recall, and f1-score for balanced dataset.
+Legitimate Email 0.99 0.97 0.98 2209
+Phishing Email 0.96 0.99 0.98 1518
+Table 4: The precision, recall, and f1-score for RoBERTa model on balanced dataset.
 Class Precision Recall F1-score Support
-Safe Email 1.00 0.97 0.98 3379
-Phishing Email 0.97 1.00 0.99 3415
-Precision: Precision which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
+Legitimate Email 1.00 0.97 0.98 2241
+Phishing Email 0.97 1.00 0.98 2288
+Precision: Precision, which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
 Precision $=\frac{TP}{TP+FP}$ (1)
@@
 Recall $=\frac{TP}{TP+FN}$ (2)
-F1-score: A balanced statistical metric that assesses a model’s efficaciousness in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall.
+F1-score: A balanced statistical metric that assesses a model’s efficacy in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall.
 It offers a thorough perspective that balances recall and precision, impacting the final score determined by the model’s projections of the positive rate and actual positive cases.
@@
 $\text{Accuracy}=\frac{\text{TN}+\text{TP}}{\text{TN}+\text{TP}+\text{FN}+\text{FP}}$ (4)
-The accuracy of a model can differ based on the various parameters, data balanced or not.
-Initially in this experiment, an imbalanced dataset was utilized to assess the model’s performance in both training and testing.
-This dataset is highly imbalanced based on the safe email and phishing email data where the number of data of safe email is quite higher than phishing email.
-To train the model with 70% imbalanced data we have used 6 epochs with a batch size of 32.
-After 6 epochs the model showed 98.90% training accuracy for this imbalanced dataset with 0.0201 training loss.
-On the other hand, we get a testing accuracy 97.50 with testing loss 0.0900 with batch size 64.
+The accuracy of a model can differ based on the various parameters, data is balanced or not.
+Initially, in this experiment, an imbalanced dataset was utilized to assess the model’s performance in training, validation, and testing.
+This dataset is highly imbalanced based on the legitimate email and phishing email data, where the number of data of legitimate emails is much higher than phishing emails.
+To train the model with 64% imbalanced data, we have used 5 epochs with a batch size of 32.
+After 5 epochs, the model showed 98.57% training accuracy for this imbalanced dataset with 0.0281 training loss.
+Then the trained model was validated with 16% unseen data and achieved 98.05% validation accuracy with 0.0513 loss.
+On the other hand, we got a testing accuracy 97.96 with testing loss 0.0560 with batch size 64.
 Fig.
 7 shows a graph of accuracy and loss with 6 epochs for an imbalanced dataset.
-Figure 6: Training vs. Testing Accuracy and Loss on Imbalanced Data
-Moreover, we also conducted this experiment with a resampled and balanced dataset.
-After preparing the balanced dataset again used 70% of the data for training purposes.
-Using this portion of data, we worked on model optimization and fine-tuning and trained the model.
-This time with balanced data the fine-tuned model showed 99.07% training accuracy with 0.0218 training loss.
-After that, we tested this trained model with 30% unseen balanced data and got 98.48% testing accuracy with 0.0447 testing loss.
-This performance showed that the performance of the model with a balanced dataset is quite better than an imbalanced dataset.
+Figure 7: Training vs. Validation vs. Testing Accuracy and Loss on Imbalanced Data
+Moreover, we also conducted this experiment with a balanced dataset.
+After preparing the balanced dataset again used 64% of the data for training purposes.
+We worked on model optimization and fine-tuning using this data portion and trained the model.
+With the balanced data, the fine-tuned model achieved 98.94% training accuracy with 0.0257 training loss, indicating exceptional performance on the training data.
+After that, the fine-tuned model was validated with 16% validation data and obtained 0.0388 validation loss with 98.68% validation accuracy, demonstrating that the model generalizes well to unseen data and is not overfitting.
+Finally, the trained model was tested with 20% unseen balanced test data and achieved 98.45% testing accuracy with 0.0511 testing loss.
+This performance showed that the performance of the model with a balanced dataset is much better than an imbalanced dataset.
 Fig.
-7 shows a graph of accuracy and loss with 6 epochs for a balanced dataset.
-The performance comparison of the model on both balanced and imbalanced dataset shows in Table 5.
-Figure 7: Training vs. Testing Accuracy and Loss on Imbalanced Data.
-Table 5: The Performance comparison of fine-tuned DistilBERT model on Imbalanced and Balanced dataset.
+8 shows a graph of accuracy and loss with 5 epochs for a balanced dataset.
+The performance comparison of the model on both balanced and imbalanced datasets is shown in Table 5.
+Figure 8: Training vs. Validation vs. Testing Accuracy and Loss on Balanced Data.
+Table 5: The Performance Comparison of Fine-Tuned RoBERTa Model on Imbalanced and Balanced Datasets.
 Dataset Mode
-Training Loss
-Training Accuracy
-Testing Loss
-Testing Accuracy
-Imbalanced
-0.0201
-98.90
-0.0900
-97.50
-Balanced
-0.0218
-99.07
-0.0447
-98.48
+Imbalanced Balanced
+Training Loss 0.0281 0.0257
+Training Accuracy 98.57 98.94
+Validation Loss 0.0513 0.0388
+Validation Accuracy 98.05 98.68
+Testing Loss 0.0560 0.0511
+Testing Accuracy 97.96 98.45
+Additionally, we employed traditional machine learning models, including Extreme Gradient Boosting (XGB), Random Forest (RF), Support Vector Machine (SVM), and Logistic Regression (LR), to evaluate their performance and compare them with our fine-tuned RoBERTa model on an imbalanced and balanced dataset.
+The generalization power and robustness of the different ML models and the RoBERTa model are shown in Table 6 and Table 7.
+Table 8 and Table 9 present a comparative performance analysis of these models.
+The overall performance analysis showed that our fine-tuned RoBERTa model outperformed the traditional ML models.
+Table 6: Precision, Recall, and F1-Score of traditional ML and RoBERTa Model for Imbalanced Dataset
+Model Class Precision Recall F1-score
+RF Legitimate Email 0.96 0.97 0.96
+Phishing Email 0.95 0.94 0.95
+XGB Legitimate Email 0.98 0.95 0.96
+Phishing Email 0.93 0.98 0.95
+SVM Legitimate Email 0.99 0.97 0.98
+Phishing Email 0.95 0.99 0.97
+LR Legitimate Email 0.98 0.96 0.97
+Phishing Email 0.94 0.97 0.96
+RoBERTa Legitimate Email 0.99 0.97 0.98
+Phishing Email 0.96 0.99 0.98
+Table 7: Precision, Recall, and F1-Score of traditional ML and RoBERTa Model for Balanced Dataset
+Model Class Precision Recall F1-score
+RF Legitimate Email 0.98 0.96 0.97
+Phishing Email 0.96 0.98 0.97
+XGB Legitimate Email 0.99 0.94 0.96
+Phishing Email 0.95 0.99 0.97
+SVM Legitimate Email 0.99 0.97 0.98
+Phishing Email 0.97 0.99 0.98
+LR Legitimate Email 0.99 0.96 0.97
+Phishing Email 0.96 0.99 0.97
+RoBERTa Legitimate Email 1.00 0.97 0.98
+Phishing Email 0.97 1.00 0.98
+Table 8: Comparative analysis of the fine-tuned RoBERTa model with traditional ML models on the imbalanced dataset.
+Model Training Loss Training Accuracy Testing Loss Testing Accuracy
+RF 0.0834 98.91 0.2241 95.65
+XGB 0.0684 98.02 0.1071 95.84
+SVM 0.0531 98.77 0.0864 97.61
+LR 0.1446 97.89 0.1707 96.46
+RoBERTa 0.0281 98.57 0.0560 97.96
+Table 9: Comparative analysis of the fine-tuned RoBERTa model with traditional ML models on the balanced dataset.
+Model Training Loss Training Accuracy Testing Loss Testing Accuracy
+RF 0.0726 99.17 0.1837 97.09
+XGB 0.0666 98.32 0.0966 96.49
+SVM 0.0422 99.01 0.0710 97.90
+LR 0.1343 98.36 0.1489 97.24
+RoBERTa 0.0257 98.94 0.0511 98.45
 Finally, we have compared our Phishing Email Detection study with some previous studies in the same field in order to assess its quality.
@@
 Table 10: Explainability of Different Studies on Phishing Email Detection
-Reference
-Datasets
-Models
-Explainability Analysis
-2023(Jamal et al., 2023)
-Email Spam Detection Dataset, PhishingEmailData
-DistilBERT, RoBERTA
-No
-2021(Alhogail and Alsabih, 2021)
-Fraud Dataset
-GCN
-No
-2019(Fang et al., 2019)
-IWSPA-AP
-THEMIS
-No
-2023(Thapa et al., 2023)
-IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl
-THEMIS, BERT
-No
-2023(Atawneh and Aljehani, 2023)
-Phishing Email Dataset, Benign Email Dataset
-BERT-LSTM
-No
-This Study Phishing Email Fine-tuned DistilBERT Yes
-The comparison showed that our research work provides transparency in this phishing email detection task.
-However, some listed research work showed higher accuracy than our proposed model.
+Reference Datasets Models Explainability Analysis
+2023(Jamal et al., 2023) Email Spam Detection Dataset, PhishingEmailData DistilBERT, RoBERTA No
+2021(Alhogail and Alsabih, 2021) Fraud Dataset GCN No
+2019(Fang et al., 2019) IWSPA-AP THEMIS No
+2023(Thapa et al., 2023) IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl THEMIS, BERT No
+2023(Atawneh and Aljehani, 2023) Phishing Email Dataset, Benign Email Dataset BERT-LSTM No
+This Study Phishing Email Fine-tuned RoBERTa Yes (LIME, Transformers Interpret, Hybrid LITA)
+The comparison showed that our research work provides transparency in the phishing email detection task.
+However, some of the listed studies showed higher accuracy than our proposed model.
 This happened because they used different datasets in their experiment.
-In this study, we only focused on the phishing email dataset but other existing works were also done this work with the spam dataset.
-Furthermore, to the best of our knowledge, none of the papers listed in Table  6 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.
+In this study, we only focused on the phishing email dataset, whereas other existing works also included the spam datasets.
+Furthermore, to the best of our knowledge, none of the papers listed in Table  10 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.
 5 Explainability Analysis
-LIME and Transformer Interpret both showed how the fine-tuned DistilBERT model arrives at its prediction on phishing email detection by approximating the model’s behaviour on every word of the email text.
-Lime and Transformer Interpret provided accurate results by answering the “Why” model predicting the selected text is a “Safe Email” or “Phishing Email”.
-Based on the explainability, those explainable techniques generate word attributions by providing positive and negative numbers on each word.
-Those numbers showed which word was more specifically reasonable for the predicted class.
-By using those numbers the explainable model visualizes the whole text by highlighting the words that scored positive and as well as negative on the prediction.
-For this explainability analysis in LIME explainer, we initialized the number of features (words) 15 for the dimensionality reduction.
-If the number of words of a selected test is very large that time the computational time will increase as well as it increases the dimensionality which also can be a reason why the systems crashed.
-Limiting this number of features makes the explanation process computationally more feasible and provides a clearer and more concise picture of which words in the case of text are influential in the model’s prediction for a particular instance.
+LIME and Transformers Interpret both demonstrated how the fine-tuned RoBERTa model arrives at its prediction for phishing email detection by approximating the model’s behavior on every word of the email text.
+These explainability methods provided accurate results by answering “Why” the model classifies a selected text as a legitimate or phishing email.
+Based on the explainability, these methods assign positive and negative importance scores to each word, indicating their contribution to the predicted class.
+These scores highlight which word was more specifically reasonable for the predicted class.
+By using those scores, the explainability models visualize the whole text by highlighting the words that scored positive or negative on the prediction.
 Fig.
-8 shows the model explainability using LIME and Table  7 and Table  8 show the positive and negative weights associated with words that help users understand phishing email detection.
-(a) Explanation for the first text
-(b) Explanation for the second text
-Figure 8: LIME Explanation for two Phishing Email Texts
-Table 7: The positive and negative weights using LIME for first text
-Positive Numbers Negative Numbers
-Word Value Word Value
-‘low’ 1.760212017947662 ‘neonate’ -0.1267345497528774
-‘price’ 1.5916451492409904 ‘3107’ -0.013956063314091143
-‘software’ 1.2586059555064724
-‘setupmefree’ 0.9966181271429526
-‘com’ 0.3984979685547945
-‘http’ 0.1447738056917774
-‘a’ 0.06584106490791478
-Table 8: The positive and negative weights using LIME for second text
-Positive Numbers Negative Numbers
-Word Value Word Value
-‘megapixel’ 0.45124215111245936 ‘a’ -0.09469945665856638
-‘get’ 0.4369880334209157
-‘free’ 0.4206519368628723
-‘digital’ 0.3721259726068145
-‘camera’ 0.2388125335048898
-‘canon’ 0.19498844421129996
-‘6’ 0.04953986530836751
-‘3’ 0.024635744735869915
-Along with LIME, we also examined the model interpretability with similar text by using Transformer Interpret.
-During this interpretation, this Transformer Interpret uses the tokenizer to split the large words into sub-words with the ‘##’ prefix.
-This happens when an unusual or non-vocabulary word is encountered by the model, which breaks it down into smaller sub-words to enable efficient processing.
-Fig.
-9 shows Transformer Interpretation for a phishing email and Table  9 and Table  10 show the word importance, with positive and negative numbers associated with each word for two phishing emails.
-(a) Explanation for first text
-(b) Explanation for second text
-Figure 9: Transformer Interpret Explanation for two Phishing Email Text
-Table 9: The positive and negative numbers using Transformer Interpret for first Phishing Email Text
-Positive Numbers Negative Numbers
-Word Value Word Value
-‘low’ 0.7698059423920272 ‘.’ -0.09537831853568463
-‘price’ 0.4465021887746993 ‘setup’ -0.037898186884516454
-‘software’ 0.10363795047092349 ‘?’ -0.04837875144898304
-‘http’ 0.05008396373925551 ‘a’ -0.212114259929307
-‘:’ 0.06724251662885174 ’=’ -0.13427331689723981
-‘/’ 0.018869178109329452 ‘310’ -0.06162690120821821
-‘neon’ 0.07499325538696087 ‘##7’ -0.013360917838605864
-‘##ate’ 0.09833698869154901 ‘.’ -0.1269640647507237
-‘me’ 0.09196861346171359
-‘free’ 0.16949987141093556
-‘com’ 0.20340796716052334
-‘/’ 0.006729002810807768
-Table 10: The positive and negative numbers using Transformer Interpret for second Phishing Email Text
-Positive Numbers Negative Numbers
-Word Value Word Value
-‘get’ 0.49733595483460524 ‘a’ -0.2688505206057748
-‘free’ 0.4645088664478948 ‘6’ -0.09532782298942542
-‘mega’ 0.18191195439136695 ‘.’ -0.08192496249808812
-‘3’ 0.014327378740528552 ‘##pi’ -0.11609630028309895
-‘##x’ 0.5855798135703232 ‘!’ -0.18838964048006548
-‘digital’ 0.021290989158880363
-‘canon’ 0.13188078942699152
-‘camera’ 0.05155301993306188
-‘>’ 0.03909299350598931
-In Fig.
-9 the highlighted text with green color shows the word importance to classify the analyzed text as Phishing.
-On the other hand, the red portions showed those words or sub-words that worked against the classified class.
-As we worked with a binary classification task here “True Label" = 1 means "Phishing Email" and when “True Label" = 0 means “Safe Email”.
-According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the two tested input texts, but they differ in how they define explanation.
-LIME mostly uses small textual modifications to evaluate the impact of the model’s prediction on the entire text, while Transformer Interpret employs an attention mechanism to examine each segment of the input sequence.
-Furthermore, in the LIME and Transformer Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email contents.
-The positive words are mainly responsible for the predicting class.
-More specifically, when those words are present in a text that time model predicts a class.
-A comprehensive difference between the two explainable AI techniques is shown in Fig.
-10 and Fig.
-11.
-The number of scoring percentages for those email texts is not the same for LIME and Transformer Interpret.
-This is only because the working procedures of those techniques are not the same.
-(a) LIME Scoring
-(b) Transformer Interpret Scoring
-Figure 10: LIME and Transformer Interpret Comparison on first text
-(a) LIME Scoring
-(b) Transformer Interpret Scoring
-Figure 11: LIME and Transformer Interpret Comparison on second text
+9 shows the LIME-based explanation for two phishing email samples, which the model also classified as phishing.
+The contributions can be seen in context by highlighting keywords in the text instances with the appropriate color.
+To better understand the explanation produced by the LIME explainer, presented bar plots of those phishing email samples by indicating the top 15 words ranked by their contribution to the classification.
+The bar length with weight indicates the strength of each word.
+Furthermore, to make the explanation relevant to the most likely result, we concentrated on the top-predicted label.
+This is especially helpful when the main goal is to comprehend the choice for the most likely class.
+(a) Explanation for the phishing sample 1
+(b) Explanation for the phishing sample 2
+Figure 9: LIME Explanation for two Phishing Email Samples
+In addition to LIME, we used Transformers Interpret to analyze model interpretability through comparable phishing examples.
+This tool examines how each token influences the model’s final prediction to generate token-level attribution scores.
+Since RoBERTa uses a BPE tokenizer, complex or rare words are broken into smaller word units, making it easier for the model to handle out-of-vocabulary terms.
+Figure 10 illustrates a visual interpretation for phishing email samples, where tokens highlighted in green contributed positively to the classification as phishing email, and tokens in red contributed negatively.
+(a) Explanation for sample 1
+(b) Explanation for sample 2
+Figure 10: Transformers Interpret Explanation for two Phishing Email Samples
+According to the analysis, both explainable approaches explain how the suggested model predicts "Phishing Email" for each of the tested input samples.
+However, those techniques differ fundamentally in how the explanations are generated.
+LIME primarily uses small textual perturbations to evaluate the impact of individual words on the model’s prediction across the entire text.
+At the same time, Transformers Interpret employs a gradient-based explanation technique to analyze the contribution of each token or sub-word to the model’s final prediction.
+Furthermore, in the LIME and Transformers Interpret explanation processes, the resulting quantity of tokens, words, or sub-words differs for the two phishing email samples.
+In both techniques, the positively contributing words are mainly responsible for the predicted class.
+More specifically, the presence of those tokens significantly increases the likelihood of the model predicting the sample as phishing.
+In our study, we found that both methods can have opposite polarities for the same words.
+For example, one explainable method finds a word to be positive and the other negative, which can lead to confusion in the mind of end-users.
+To address this issue, we developed our hybrid approach, LITA, for our analysis.
+By integrating the complementary strengths of the two methods, we achieve a more balanced explanation.
+The LITA also generated explanations for two phishing samples shown in Fig.
+11; and Table 11 and Table 12 present the top 15 words ranked by their LIME scores for each sample, along with the corresponding scores from Transformers Interpret and the hybrid LITA approach.
+These tables present quantitative evidence that highlights interpretability differences between both individual methods and illuminates how LITA reconciles aligned and conflicting attributions to improve clarity and trust.
+(a) Explanation for the phishing sample 1
+(b) Explanation for the phishing sample 2
+Figure 11: LITA Explanation for two Phishing Email Samples
+Table 11: Word-Level Explainability Analysis for Phishing Sample 1 Using LIME, Transformers Interpret, and Hybrid Scoring
+Word LIME Score LIME (Norm) Score Transformer Score Transformer (Norm) Score Hybrid Score
+freida 0.35498 1.00000 0.10828 0.45581 0.72791
+verify 0.32802 0.91895 -0.01591 -0.32235 0.29830
+senior -0.31029 -1.00000 0.00691 -0.17936 -0.58968
+confirm 0.29308 0.81391 -0.06232 -0.61315 0.10038
+loan 0.27749 0.76704 0.02076 -0.09258 0.33723
+here 0.25293 0.69321 0.00004 -0.22241 0.23540
+webber 0.25283 0.69291 0.16803 0.83020 0.76155
+applicant -0.24042 -0.78995 -0.12406 -1.00000 -0.89497
+application -0.21336 -0.70860 -0.08878 -0.77894 -0.74377
+a -0.19875 -0.66468 0.02981 -0.03587 -0.35027
+financial 0.19619 0.52263 0.08843 0.33143 0.42703
+000 0.18949 0.50249 0.10274 0.42110 0.46179
+association 0.18149 0.47844 0.19513 1.00000 0.73922
+your 0.17999 0.47393 -0.00322 -0.24283 0.11555
+manager -0.17817 -0.60281 -0.01086 -0.29070 -0.44676
+Table 12: Word-Level Explainability Analysis for Phishing Sample 2 Using LIME, Transformers Interpret, and Hybrid Scoring
+Word LIME Score LIME (Norm) Score Transformer Score Transformer (Norm) Score Hybrid Score
+free 0.18759 1.00000 0.13196 0.39400 0.69700
+perfect 0.14421 0.64174 0.15117 0.46549 0.55361
+removed 0.10480 0.31627 0.18932 0.60745 0.46186
+want 0.09803 0.26035 0.20522 0.66662 0.46349
+remove 0.08880 0.18413 0.13166 0.39289 0.28851
+special 0.08735 0.17215 0.13495 0.40513 0.28864
+no 0.06542 -0.00896 -0.21209 -0.88628 -0.44762
+type 0.05757 -0.07379 0.01917 -0.02571 -0.04975
+somebody 0.05757 -0.07379 0.00061 -0.09478 -0.08429
+the -0.05458 -1.00000 0.03316 0.02635 -0.48683
+gift 0.05245 -0.11608 0.29481 1.00000 0.44196
+be 0.04975 -0.13837 0.00989 -0.06025 -0.09931
+further 0.04455 -0.18132 -0.24265 -1.00000 -0.59066
+catalog -0.04401 -0.91271 -0.21059 -0.88070 -0.89670
+please -0.03814 -0.86423 0.02428 -0.00670 -0.43546
 6 Discussion
-The results of both an imbalanced and balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned DistilBERT model.
-The fine-tuned version of the DistilBERT model showed 98.90% training and 97.50% testing accuracy on an imbalanced dataset whereas this model showed 99.07% training and 98.48% testing accuracy on the balanced dataset.
-Although these results show that the model’s performance on a balanced dataset is quite better than an imbalanced dataset, the difference is not so high, approximately 1%.
+The results of both an imbalanced and a balanced dataset depict an enhancement in the performance of the proposed transformer-based fine-tuned RoBERTa model.
+The fine-tuned version of the RoBERTa model showed 98.57% training and 97.96% testing accuracy on an imbalanced dataset, whereas this model showed 98.94% training and 98.45% testing accuracy on the balanced dataset.
+Although these results show that the model’s performance on a balanced dataset is quite better than on an imbalanced dataset, the difference is not so high, approximately 0.49%.
 This demonstrated that the fine-tuned version performs well in an imbalanced dataset.
 Additionally, the performance shows that the optimized model performed better than other previous studies.
-Besides, this research work also answered the “How was the model performed?" and “Why was it performed?" using XAI, motivated from recently published book by Sarker (Sarker, 2024).
-Two well-known Explainable and interpretable techniques are utilized for the model understanding.
-It is more difficult to understand how DistilBERT makes its predictions because, as we know, it is a complex model.
+Additionally, this research work also addressed the questions “How was the LLM model performed?" and “Why was it performed?" using XAI, as motivated by the recently published paper by Sarker (Sarker, 2024).
+The author highlighted the potentiality of the LLMs by addressing the technical, ethical, and societal challenges, focusing on fairness, transparency, transparency, explainability, and trust.
+By summarizing and understanding this potential research scope with direction, two well-known explainable and interpretable techniques are utilized in our research to understand the model’s decision.
+It is more difficult to understand how RoBERTa makes its predictions because, as we know, it is a complex model.
 These AI techniques aided in gaining an understanding of the model’s decision-making process.
-Both LIME and Transformer Interpret highlighted the areas of the text that are more responsible for the classification of the phishing of safe email.
+Both LIME and Transformers Interpret highlighted the areas of the text that are more responsible for the classification of phishing or legitimate emails.
 This highlights areas worked with words or sub-words, making the decision more understandable for humans, and users can also comprehend it easily.
-Mainly, those techniques calculated each word or sub-words numeric value or weight where some words indicate positive value and some negative.
+Mainly, those techniques calculated each word or sub-words numeric value or weight, where some words indicate positive value and some negative.
 Words with calculated higher positive values are more important for classifying the predicted class, while those with more negative values do not support the predicted class.
 This experimental analysis showed the transparency of the model in this phishing email classification task.
+Although those methods show almost similar explanations, but in some word cases their explanation not agreed due to their different functionality.
+But this can create serious confusion which one solved in this work, with LITA.
 7 Conclusion
-This study has investigated the detection of phishing emails by utilizing a Large Language Model.
-This also has shown how effective it is to use state-of-the-art natural language processing methods to improve the precision and effectiveness of phishing email identification.
-The overall experiment was done with both a balanced and imbalanced dataset.
-We have obtained promising results in differentiating between phishing and valid emails by utilizing DistilBERT’s contextual language understanding technique.
-Mainly, to achieve this good performance the model is optimized by changing some important parameters such as batch size, optimization function, learning rate, etc, and also fine-tuning has been done.
-This fine-tuned version of DistilBERT obtained 98.48% testing accuracy in this phishing email classification task.
-Furthermore, LIME and Transformer Interpret techniques have been employed to see the explanation of the model and how it makes decisions.
-The decision-making process is now more visible and understandable for users because of the usage of explainable AI, which promotes comprehension and trust.
-The overall model explainability describes the robustness of the proposed model.
-The field of Large Language Models has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers.
+This study explores the detection of phishing emails using a transformer-based language model.
+It also demonstrates the effectiveness of employing state-of-the-art natural language processing methods to improve the accuracy and efficiency of identifying phishing emails.
+The overall experiment was conducted with both balanced and imbalanced datasets.
+We achieved promising results in distinguishing between phishing and legitimate emails by utilizing RoBERTa’s contextual language understanding capabilities.
+To accomplish this, the model was optimized by adjusting key parameters such as batch size, optimization function, learning rate, and through fine-tuning.
+This fine-tuned version of RoBERTa obtained a 98.45% testing accuracy in the phishing email classification task.
+Additionally, to improve interpretability and foster greater trust in the model’s explanations, we introduced LITA, a hybrid explanation approach that combines LIME and Transformers Interpret.
+This makes the decision-making process more transparent and understandable for users, promoting comprehension and trust.
+Overall, the model’s explainability reflects the robustness of the proposed system.
+The field of LLMs has experienced quick and dynamic development due to the substantial investment made by both businesses and consumers.
 New open-source models are being produced almost every day, demonstrating the ongoing development and innovation in this field.
-In the future, we plan to expand the scope of our research by experimenting with a wider variety of models and datasets.
-Furthermore, we intend to investigate the use of different variants of BERT-based models for the identification of different language phishing emails.
-Moreover, deep analysis of the model will be conducted to ensure its robustness by enhancing the explainable AI techniques.
+In future work, we plan to explore various LLMs, including LLaMA, Gemini, Falcon, Mistral, and others, to assess their effectiveness in phishing email detection.
+We will extend our research to multilingual datasets to improve our model’s performance across different languages and cultural contexts.
+To further enhance the robustness and accuracy of our algorithms, we also plan to create a new dataset comprising phishing emails.
+Additionally, we aim to integrate our models into real-world cybersecurity systems, partnering with industry players to evaluate their real-time effectiveness and refine them based on practical feedback.
+Furthermore, we will conduct more in-depth analysis of the models’ decision-making processes to ensure robustness by advancing explainable AI techniques.
 […]
 Data Availability Statement
-The dataset that is used in this study is available at Kaggle (Online Link):  https://www.kaggle.com/datasets/subhajournal/phishingemails.
+The research dataset which one is used in this study is available at Kaggle:  https://www.kaggle.com/datasets/subhajournal/phishingemails.